OpenStudy (anonymous):

A true-false test consists of 20 questions, each of which has one correct answer: true, or false. One point is awarded for every correct answer, but one point is taken off for each wrong answer. Suppose a student answers every question by guessing at random, independently of other questions.

OpenStudy (anonymous):

E(S)=0 SE(S)= 4.472 P(S=0)= 0.1761

OpenStudy (anonymous):

Let Xi=1 if the student is right on the i-th question, and 0 if she is not right. Let Mi be the student's mark on the i-th question. Then Mi=2Xi−1. The test score S is then given by S=M1+M2+⋯+M20. It is standard that the standard error of Xi is √(1/2)(1−1/2)=1/2. So the standard error of 2Xi−1 is (2)(1/2)=1. The standard error of the sum M1+M2+⋯+M20 is therefore (√20)(1). Remark: when you divided the (not quite right) standard error per question by √20, you were finding the standard error of the average mark per question, not the standard error of the total mark. We could also find the SE of each Mi directly. The variance of Mi is E(M2i)−(μi)2, where μi is the mean of Mi. But the mean of Mi is 0. And M2i=1 always, so E(M2i)=1. It follows that the variance of Mi is 1, and therefore the SE of Mi is 1. The Mi are assumed independent. The total mark is the sum of the Mi, so has variance (20)(1), and therefore standard deviation √20.

OpenStudy (kenljw):

P(C) = .5, P(E) = .5 P(C) - P(E) = 0 This indicates a mean of 0 Use binomial distribution to determine probability between -20 to 20